find the slope of the line parallel to the line -2x + y = 3
Answer by mducky2(62)

Firstly, you must solve for x.
-2x + y = 3
-2x + 2x + y = 3 + 2x
y = 3 + 2x

Any line parallel to this line has to have the same slope, so just find the slope of this equation. After you solve for y, the slope of any linear equation is the multiplier of x. In this case, y = 3 + 2x, the number 2 is the slope because its the multiplier of x. Therefore, the slope for any parallel line is the same: 2.